### v1

@@ 1.3 Mathematical model of a $\pi^{2}$ neuromorphic architecture
where $W^{n}$ is the synaptic weight matrix connecting layer n1{n-1} to layer $n$, $\sigma(\cdot)$ denotes the activation function applied at the output of every layer and $\odot$ represents the $\pi^{2}$ operations based on the CBS and ATS primitives described in Fig. 2. Note that for a conventional deep neural networks, the $\odot$ in equation 4 are based on the standard multiply-and-accumulate (MAC) operations interleaved with nonlinear transformations.
$\pi^{2}$-NN use synaptic weights to introduce programmable delays in the arrival times of input events. A differential margin-propagation proposed in [23],[29] is used to approximate inner products and is applied to Eq. 1. Accordingly, the weights and inputs of the $\pi^{2}$ network are represented in a differential formulation. The real-values d-dimensional input X is encoded into differential event spike times and fed as input to $\pi^{2}-NN$ as follows:
@@
The synaptic weights (W) are encoded as differential synaptic delays represented as $W^{+}=[B+W]_{+},W^{-}=[B-W]_{+}$ for all layers. Here, A and B are arbitrary constants, and $_{+}=\max(0,.)$ is a ReLU operation that ensures that the information represented in time is non-negative. The output of a $j^{th}$ neuron in a $\pi^{2}$-NN is two spikes/events denoted by their respective time of occurrences $T_{j}^{+}$ and $T_{j}^{-}$. These occurences are computed as follows:
$T^{+}_{j}=\frac{M}{K}+\frac{1}{K}\sum_{k=1}^{K}t_{[k],j}^{+}\ \text{and}\ \ T^{-}_{j}=\frac{M}{K}+\frac{1}{K}\sum_{k=1}^{K}t_{[k],j}^{-}\quad\text{where}$ (6)
@@
The terms $t_{[1]},t_{[2]},\ldots,t_{[K]}$ represent the first $K$ earliest arrival times among the inputs defined in the sets. As described in [29], the difference between the positive and negative differential spike times provides an approximation of a dot product operation. To obtain a single-ended output, the differential spike times are combined as follows, and a nonlinear activation function (ReLu) $\sigma(.)$ is subsequently applied:
$)T_{j}=\sigma\ (\alpha(T^{-}_{j}-T^{+}_{j}$)) (9)
Here, $\alpha$ is a hyperparameter that helps in faster convergence during training. The single-ended output in Eq.9 is subsequently re-encoded into a differential form, consistent with the input encoding described in Eq.5, and propagated to the next layer for further processing. Such a mapping enables the embedding of event-driven computation within conventional Ethernet-based communication protocols. This approach opens the door to leveraging ongoing advancements in switching and routing hardware to efficiently scale neuromorphic computing architectures.
@@ 2 Results
Figure 4: A) Scatter plot portraying the relationship between weights of the $\pi^{2}$ and MLP networks (784x50x10 architecture) trained on the MNIST dataset. We chose this network due to the memory constraints of running the simulator on our system. When the trained MLP weights are directly mapped to the $\pi^{2}$ network weights, there is a 1-1 relationship. After retraining the $\pi^{2}$ network, a high correlation still exists. B) t-SNE plot of the hidden layer representations of the ten classes extracted from the trained MAC-based MLP network. Distinct clusters are obtained from different classes of the trained network. C) When the $\pi^{2}$ network is initialized with the trained MLP weights, the separation between the clusters still exists as it tries to approximate the MAC operation. However, an approximation error exists. This error propagates through the next consecutive layers, making the classification accuracy drop from 97.2% to 88.7%. Each color represents a distinct class in these plots. D,E) The $\pi^{2}$ network has to be trained further to compensate for the drop in classification accuracy. The effect of the scaling parameter $\alpha$ on the test accuracy and training loss is portrayed here. Higher $\alpha$ helps in improving the separation between the classes to achieve robust classification. F) The 3-layer $\pi^{2}$ network with 50 nodes ($\pi^{2}(50)$) is retrained to achieve a classification accuracy of 97.34%. The weights of the trained $\pi^{2}$ network are quantized to 3 bits ($Q\pi^{2}$) and simulated using Ethernet switches on OMNET++ software. The simulator supports 3-bit PCP codes and a network with 50 nodes, so we simulate accordingly. The simulation results (sim) exactly match the quantized networks’ outputs to achieve a test accuracy of 96.67%. By increasing the number of hidden nodes to 800, we can achieve the baseline accuracy reported in [30] ($\pi^{2}(800)$).
@@
Figure 5: A,B) The spatio-temporal patterns emerging from the layers (input (Layer 1), hidden (Layer 2) and output layer (Layer 3)) of the trained $\pi^{2}$ network are illustrated as a raster plot for different configurations of K. The blue and green dots represent the differential event generation times of the nodes in a layer. The population response to the input digit “4” is shown for (A) $K=[1,16]$ and (B) $K=[10,16]$ for the hidden and output layers, respectively. C,D) The corresponding evolution of neuronal activity for the same input is depicted for (C) $K=[1,16]$ and (D) $K=[10,16]$ (hidden and output layers). It can be observed that reducing the value of K to 1 leads to a computationally sparse (in terms of neuron and synaptic activity) and faster computation. Differential event activity (based on Eq. 5,9 is selectively traced for classes “4” and “1” in these plots.
@@
We first explore the functional equivalence between conventional neural networks using (multiply-accumulate, and non-linear activation operations) and $\pi^{2}$-NNs for a small-scale problem where we can analyze the intrinsic dynamics and learning in more detail. We consider a standard 3-layer multi-layer perceptron (MLP) with a 784-50-10 architecture and ReLU activation at the hidden layer, trained on the MNIST dataset. This particular configuration was selected because it could be executed using the OMNeT++ network simulator [31]. The OMNET++ simulator is widely used in the networking literature [32, 27, 33, 34] to benchmark different packet-routing and congestion control algorithms. As a result OMNET++ simulator supports many of the existing Ethernet protocol standards. However, due to the simulator memory constraints, this network architecture was chosen to simulate a model with fewer than 1,000 nodes. A comparable or smaller scale node count is also adopted for scalability studies using the OMNET++ simulator [27],[35]. For this architecture, the MLP achieves a test accuracy of 97.2%. The learned weights are then directly mapped to an equivalent $\pi^{2}$-NN, enabling evaluation of the model’s performance in the time domain. For specific values of the scaling parameter $\alpha=[30,30]$ and $K=[140,16]$ for the hidden and output layer, respectively, the $\pi^{2}$-NN achieves a test accuracy of 88.7%. Despite the performance gap, both models exhibit strong functional resemblance in their learned representations. This similarity is visualized in Fig. 4A,B,C using t-SNE projections of the hidden-layer activations from the MLP and the corresponding $\pi^{2}$-NN. A formal derivation of their equivalence is provided in the Methods section. Furthermore, we show that retraining the $\pi^{2}$-NN can recover the performance of the original MLP, effectively closing the performance gap introduced by direct weight transfer (Fig. 4D,E). Finally, we note the critical role of the scaling parameter $\alpha$ in training. Larger values of $\alpha$ enhance class separability and accelerate convergence during training. Fig.4D,E illustrates this behavior by comparing training trajectories for different values of $\alpha$. Furthermore, in contrast to conventional neural networks, information in $\pi^{2}$-NNs is encoded spatiotemporally, leveraging both the spatial structure of the network and the temporal dynamics of event-driven computation. This encoding paradigm enables $\pi^{2}$-NNs to exhibit rich and diverse combinatorial representational capabilities, as discussed in [36]. The emergent spatiotemporal representations across successive layers of the $\pi^{2}$-NN and the impact of the parameter K on the computational sparsity and latency of the networks are illustrated in Fig. 5. The spatio-temporal activity observed for inputs of different classes is presented in the supplementary section.
@@
To demonstrate the $\pi^{2}$-NN architecture can be mapped onto current packet-switching protocols, we implemented this network using Ethernet switches on the OMNeT++ network simulation platform [32, 27]. The simulator implements the core traffic shaping protocols as discussed in the previous section. The simulation results exactly match the software outputs of the 3-bit quantized network, as shown in Fig. 4F. The spatio-temporal patterns (similar to Fig. 5) extracted from the simulator are presented in the supplementary section. Due to memory limitations, simulation results are presented only for a small 3-layer $\pi^{2}$-NN comprising slightly fewer than 1,000 nodes. Notably, the differential nature of the $\pi^{2}$ algorithm effectively doubles the number of traffic flows, further increasing the simulation complexity. However, the accuracy of the network can be improved by increasing the number of hidden nodes to 800 (Fig. 4F), and we can match the baseline accuracy reported in [30]. Fig. 4F highlights that a 3-bit quantized $\pi^{2}$-NN can be mapped to the simulator without any drop in accuracy. So, for larger-scale $\pi^{2}$-NNs, scalability analysis is performed in software, and the corresponding results are reported based solely on software-level simulations in the subsequent experiments.
@@
Figure 6: Evaluation of $\pi^{2}$ Network Performance and Robustness. A) The test accuracy obtained on different $\pi^{2}$-NN architectures trained on standard image classification datasets - MNIST, FMNIST, CIFAR10, CIFAR100 is reported. Average classification accuracy over 10 runs is reported here. B) Test accuracy obtained on a $\pi^{2}$ based 3-layer network trained on the MNIST dataset with varying hidden node dimensions. The blue bar represents the accuracies obtained when K is constrained to 1 at the hidden layer. The green bar represents the accuracies obtained when K is constrained to 1 at the hidden and output layers. C,D) Layer-wise configuration of $K$ in $\pi^{2}$ based LeNet-5 and ResNet9 models trained on FMNIST and CIFAR-10 respectively, showing the trade-off between computational sparsity and classification accuracy. E) Effect of weight quantization on accuracy with and without fine-tuning. F) Accuracy as a function of the standard deviation (SD) of random noise values added to each spike time in the network (spiking time jitter). On fine-tuning for 10 epochs, the performance drop due to quantization and jitter can be significantly improved.
@@
Figure 7: Evaluation of $\pi^{2}$ Network’s Robustness. A) Effect of weight quantization on accuracy with and without fine-tuning. B) Accuracy as a function of the random packet/event drop probability applied independently at every layer. C) Accuracy as a function of the standard deviation (SD) of random noise values added to each spike time in the network (spiking time jitter). On fine-tuning for 10 epochs, the performance drop due to quantization, jitter, and random packet drops can be significantly improved. The average accuracy across 10 trials is reported here.
@@
We perform similar experiments on the F-MNIST dataset trained on the LeNet-5 network architecture comprising 60K parameters. We train the traditional MAC-based network and achieve an accuracy of 90.47%. On porting the weights to an equivalent $\pi^{2}$-NN network with appropriately chosen values of $K$ and $\alpha$, we achieve an initial test accuracy of 85.23%. Re-training the $\pi^{2}$-NN improves the accuracy to 90.7%. Training a $\pi^{2}$-NN network from scratch with different random initializations leads to a 1% accuracy drop on average. For comparison, [30] reports a test accuracy of 90.94% using a TTFS (Time-To-First-Spike) spiking neural network (SNN) variant of LeNet, while [37] reports 90.1% using an SNN-based LeNet model traned on the FMNIST dataset.
@@
Nevertheless, upon further training using the initialized weights, the $\pi^{2}$-ResNet-9 achieves a competitive maximum accuracy of 94.23% on CIFAR-10. Prior works on ANN-based models report accuracies of 94.29% and 90.8% in [38] and [39], respectively. A custom 9-layer residual convolutional network, DavidNet, achieves 94.08% accuracy [40], while a spiking ResNet-18 model (which has around 11.2M parameters) demonstrates 94.25% accuracy in [41]. Notably, 94% is also considered the benchmark for human-level performance on CIFAR-10 [42].
@@
For the CIFAR-100 dataset, $\pi^{2}$-based ResNet-9 model achieves a maximum test accuracy of 75.5%. In comparison, previous studies report accuracies of 75.41% and 68.25% on ResNet-9-based DNN models, in [43] and [44], respectively. In [45], performance is further improved through model ensembling, resulting in an accuracy of 77%. However, the accuracy achieved using a single ResNet-9 model in that study is 72.21%. Further, researchers report an accuracy of 74.24 % on ResNet-18-based SNNs in [41]. These comparative results are summarized in Fig. 6A. Training $\pi^{2}$-NNs on GPUs is both memory-intensive and computationally demanding due to their differential formulation and the nature of their core operations. While GPUs are highly optimized for parallel matrix–vector multiplications, $\pi^{2}$-NN training requires sorting and event-driven operations that are not inherently GPU-friendly [46]. As a result, we were unable to demonstrate results for significantly larger network architectures. All our experiments were conducted on a single 40GB NVIDIA A100 GPU. Training a $\pi^{2}$-ResNet-9 network on the CIFAR datasets with a batch size of 16 required approximately 34GB of GPU memory, thus constraining the training of larger scale networks. Further details of the training configuration are provided in the Methods section 4.3.
@@
Fine-tuning for hardware: One critical architectural lever for optimizing hardware efficiency is the choice of the parameter K. Reducing the value of K lowers the computational latency and improves computational sparsity (as fewer inputs need to be processed, as portrayed in 5C,D), reduces hardware complexity and minimizes the buffer and queue size requirements. Specifically, a simplified processor design can be implemented by setting K=1 in Eq. 1. The $\pi^{2}$-Rx node generates an event immediately upon receiving the first input event, reducing the neuron model to a basic processing element—a switch. This kind of paradigm evaluates the interconnects’ computational capabilities alone. While this approach significantly improves throughput by minimizing computational complexity to that of a switching operation (Methods section 4.7) and reducing latency, it comes at the cost of reduced accuracy (Fig. 6 B,C,D). Further, the $K=1$ networks also eliminate the memory access overhead of fetching the pre-computed values (membrane potential) from memory, thus enhancing the energy efficiency of $\pi^{2}$ networks [47, 48]. However, it is important to note that the performance of $\pi^{2}$-NNs, as observed in software simulations, may degrade in real hardware due to physical limitations and non-idealities of the interconnection networks. Therefore, the network must be designed to be robust against hardware-level effects such as quantization, temporal noise (jitter) and packet drops.
@@
Analyzing the tradeoffs between K and computational accuracy: We evaluate the computational capabilities of the interconnects using the $K=1$ paradigm within a 3-layer fully connected network trained on the MNIST dataset. Initially, $K$ is fixed at 1 in the hidden layer, while the output layer is allowed to use $K>1$ during training. Subsequently, we constrain $K$ to 1 in the output layer as well, to assess the network’s performance under fully constrained conditions. The results are summarized in Fig. 6B. A baseline accuracy of 97.7% was previously reported for a 3-layer MLP with 100 hidden neurons [23]. Applying the $K=1$ constraint in the hidden layer of the same architecture results in a 2% drop in accuracy. However, increasing the hidden layer dimension to 500 restores the original performance. The trade-off in accuracy when K is constrained to 1 in the output layer as well is also reported.
@@
We extend a similar experiment to the $\pi^{2}$-NN-based LeNet-5 architecture trained on the FMNIST dataset. As illustrated in Fig. 6C, imposing the $K=1$ constraint on the first four layers of the network leads to an approximate 10% reduction in test accuracy relative to the baseline. To mitigate this drop, we relax the constraint by setting $K=10$ across all layers. This results in a more modest accuracy degradation of approximately 3%, highlighting the trade-off between computational sparsity and model performance.

### v3

@@ 1.3 Mathematical model of a $\pi^{2}$ neuromorphic architecture
where $W^{n}$ is the synaptic weight matrix connecting layer n1{n-1} to layer $n$, $\sigma(\cdot)$ denotes the activation function applied at the output of every layer, and $\odot$ represents the $\pi^{2}$ operations based on the CBS and ATS primitives described in Fig. 2. Note that for a conventional deep neural network, the $\odot$ in equation 4 is based on the standard multiply-and-accumulate (MAC) operations interleaved with nonlinear transformations.
$\pi^{2}$-NN uses synaptic weights to introduce programmable delays in the arrival times of input events. A differential margin-propagation proposed in [temp],[mulp_mp] is used to approximate inner products and is applied to Eq. 1. Accordingly, the weights and inputs of the $\pi^{2}$ network are represented in a differential formulation. The real-values d-dimensional input X is encoded into differential event spike times and fed as input to $\pi^{2}-NN$ as follows:
@@
The synaptic weights (W) are encoded as differential synaptic delays represented as $W^{+}=[B+W]_{+},W^{-}=[B-W]_{+}$ for all layers. Here, A and B are arbitrary constants, and $_{+}=\max(0,.)$ is a ReLU operation that ensures that the information represented in time is non-negative.
The output of a $j^{th}$ neuron in a $\pi^{2}$-NN is two spikes/events denoted by their respective time of occurrences $T_{j}^{+}$ and $T_{j}^{-}$. These occurrences are computed as follows:
$\tilde{T}^{+}_{j}=\frac{M}{K}+\frac{1}{K}\sum_{k=1}^{K}t_{[k],j}^{+}\ \text{and}\ \ \tilde{T}^{-}_{j}=\frac{M}{K}+\frac{1}{K}\sum_{k=1}^{K}t_{[k],j}^{-}\quad\text{where}$ (6)
@@
The terms $t_{[1]},t_{[2]},\ldots,t_{[K]}$ represent the first $K$ earliest arrival times among the inputs defined in the sets. As described in [mulp_mp], the difference between the positive and negative differential spike times provides an approximation of a dot product operation. To obtain a single-ended output, the differential spike times are combined as follows, and a nonlinear activation function (ReLu) $\sigma(.)$ is subsequently applied:
$)T_{j}=\sigma\ (\alpha(\tilde{T}^{-}_{j}-\tilde{T}^{+}_{j}$)) (9)
$T^{+}_{j}=V+T_{j},T^{-}_{j}=V-T_{j}$ (10)
Here, $\alpha$ is a hyperparameter that helps in faster convergence during training. The single-ended output in Eq.9 is subsequently re-encoded into a differential form (Eq. 10), consistent with the input encoding described in Eq.5, and propagated to the next layer for further processing. $V$ is a constant chosen such that causality in the time domain is maintained between the input and output events. Such a mapping enables the embedding of event-driven computation within conventional Ethernet-based communication protocols. This approach opens the door to leveraging ongoing advancements in switching and routing hardware to efficiently scale neuromorphic computing architectures. Further details on the mathematical framework required for training $\pi^{2}$ neural networks using gradient-based methods are presented in the supplementary section 2.1,2.2.
@@ 2 Results
Figure 4: A) Scatter plot portraying the relationship between weights of the $\pi^{2}$ and MLP networks (784x50x10 architecture) trained on the MNIST dataset. We chose this network due to the memory constraints of running the simulator on our system. When the trained MLP weights are directly mapped to the $\pi^{2}$ network weights, there is a 1-1 relationship. After retraining the $\pi^{2}$ network, a high correlation still exists. B) t-SNE plot of the hidden layer representations of the ten classes extracted from the trained MAC-based MLP network. Distinct clusters are obtained from different classes of the trained network. C) When the $\pi^{2}$ network is initialized with the trained MLP weights, the separation between the clusters still exists as it tries to approximate the MAC operation. However, an approximation error exists. This error propagates through the next consecutive layers, making the classification accuracy drop from 97.2% to 88.7%. Each color represents a distinct class in these plots. D,E) The $\pi^{2}$ network has to be trained further to compensate for the drop in classification accuracy. The effect of the scaling parameter $\alpha$ on the test accuracy and training loss is portrayed here. Higher $\alpha$ helps in improving the separation between the classes to achieve robust classification. F) The 3-layer $\pi^{2}$ network with 50 nodes ($\pi^{2}(50)$) is retrained to achieve a classification accuracy of 97.34%. The weights of the trained $\pi^{2}$ network are quantized to 3 bits ($Q\pi^{2}$) and simulated using Ethernet switches on OMNET++ software. The simulator supports 3-bit PCP codes and a network with 50 nodes, so we simulate accordingly. The simulation results (sim) exactly match the quantized networks’ outputs to achieve a test accuracy of 96.67%. By increasing the number of hidden nodes to 800, we can achieve the baseline accuracy reported in [stanojevic2024] ($\pi^{2}(800)$).
@@
Figure 5: A,B) The spatio-temporal patterns emerging from the layers (input (Layer 1), hidden (Layer 2), and output layer (Layer 3)) of the trained $\pi^{2}$ network are illustrated as a raster plot for different configurations of K. The blue and green dots represent the differential event generation times of the nodes in a layer. The population response to the input digit “4” is shown for (A) $K=[1,16]$ and (B) $K=[10,16]$ for the hidden and output layers, respectively. C,D) The corresponding evolution of neuronal activity for the same input is depicted for (C) $K=[1,16]$ and (D) $K=[10,16]$ (hidden and output layers). It can be observed that reducing the value of K to 1 leads to a computationally sparse (in terms of neuron and synaptic activity) and faster computation. Differential event activity (based on Eq. 5,9 is selectively traced for classes “4” and “1” in these plots.
@@
We first explore the functional equivalence between conventional neural networks using (multiply-accumulate, and non-linear activation operations) and $\pi^{2}$-NNs for a small-scale problem where we can analyze the intrinsic dynamics and learning in more detail. We consider a standard 3-layer multi-layer perceptron (MLP) with a 784-50-10 architecture and ReLU activation at the hidden layer, trained on the MNIST dataset. This particular configuration was selected because it could be executed using the OMNeT++ network simulator [omnet1]. The OMNET++ simulator is widely used in the networking literature [omnet_tsn, nesting, omnet2, atscbs2, fault] to benchmark different packet-routing and congestion control algorithms. As a result, the OMNET++ simulator supports many of the existing Ethernet protocol standards. Several studies show that OMNeT++/INET can be aligned closely with real TSN hardware behavior [tsn_hw_1, tsn_hw_2]. These studies demonstrate that the traffic-shaping primitives, such as the CBS and Time-aware shaping (TAS) protocols, can be reliably assessed in simulation and that the differences between simulated and hardware behavior are well-characterized and quantifiable. However, due to the simulator memory constraints, this network architecture was chosen to simulate a model with fewer than 1,000 nodes. A comparable or smaller scale node count is also adopted for scalability studies using the OMNET++ simulator in [nesting, omn_sc1, omn_sc2]. For this architecture, the MLP achieves a test accuracy of 97.2%. The learned weights are then directly mapped to an equivalent $\pi^{2}$-NN, enabling evaluation of the model’s performance in the time domain. For specific values of the scaling parameter $\alpha=[30,30]$ and $K=[140,16]$ for the hidden and output layer, respectively, the $\pi^{2}$-NN achieves a test accuracy of 88.7%. Despite the performance gap, both models exhibit strong functional resemblance in their learned representations. This similarity is visualized in Fig. 4A,B,C using t-SNE projections of the hidden-layer activations from the MLP and the corresponding $\pi^{2}$-NN. A formal derivation of their equivalence is provided in the Methods section. Furthermore, we show that retraining the $\pi^{2}$-NN can recover the performance of the original MLP, effectively closing the performance gap introduced by direct weight transfer (Fig. 4D,E). Finally, we note the critical role of the scaling parameter $\alpha$ in training. Larger values of $\alpha$ enhance class separability and accelerate convergence during training. Fig.4D,E illustrates this behavior by comparing training trajectories for different values of $\alpha$. Furthermore, in contrast to conventional neural networks, information in $\pi^{2}$-NNs is encoded spatiotemporally, leveraging both the spatial structure of the network and the temporal dynamics of event-driven computation. This encoding paradigm enables $\pi^{2}$-NNs to exhibit rich and diverse combinatorial representational capabilities, as discussed in [poly]. The emergent spatiotemporal representations across successive layers of the $\pi^{2}$-NN and the impact of the parameter K on the computational sparsity and latency of the networks are illustrated in Fig. 5. The spatio-temporal activity observed for inputs of different classes is presented in the supplementary section 4(Fig. S3).
@@
To demonstrate the $\pi^{2}$-NN architecture can be mapped onto current packet-switching protocols, we implemented this network using Ethernet switches on the OMNeT++ network simulation platform [omnet_tsn, nesting]. The simulator implements the core traffic shaping protocols as discussed in the previous section. The simulation results exactly match the software outputs of the 3-bit quantized network, as shown in Fig. 4F. The spatio-temporal patterns (similar to Fig. 5) extracted from the simulator are presented in the supplementary section 4 (Fig. S9). Due to memory limitations, simulation results are presented only for a small 3-layer $\pi^{2}$-NN comprising slightly fewer than 1,000 nodes. Notably, the differential nature of the $\pi^{2}$ algorithm effectively doubles the number of traffic flows, further increasing the simulation complexity. However, the accuracy of the network can be improved by increasing the number of hidden nodes to 800 (Fig. 4F), and we can match the baseline accuracy reported in [stanojevic2024]. Fig. 4F highlights that a 3-bit quantized $\pi^{2}$-NN can be mapped to the simulator without any drop in accuracy. So, for larger-scale $\pi^{2}$-NNs, scalability analysis is performed in software, and the corresponding results are reported based solely on software-level simulations in the subsequent experiments. Note that our OMNeT++ results should be viewed as functional validation of the time-domain dynamics of $\pi^{2}$ rather than hardware-level performance claims.
@@
Figure 6: Evaluation of $\pi^{2}$ Network Performance and Robustness. A) The test accuracy obtained on different $\pi^{2}$-NN architectures trained on standard image classification datasets - MNIST, FMNIST, CIFAR10, and CIFAR100 is reported. Average classification accuracy over 10 runs is reported here. B) Test accuracy obtained on a $\pi^{2}$ based 3-layer network trained on the MNIST dataset with varying hidden node dimensions. The blue bar represents the accuracy obtained when K is constrained to 1 at the hidden layer. The green bar represents the accuracy obtained when K is constrained to 1 at the hidden and output layers. C,D) Layer-wise configuration of $K$ in $\pi^{2}$ based LeNet-5 and ResNet9 models trained on FMNIST and CIFAR-10, respectively, showing the trade-off between computational sparsity and classification accuracy.
@@
Figure 7: Evaluation of $\pi^{2}$ Network’s Robustness. A) Effect of weight quantization on accuracy with and without fine-tuning. B) Accuracy as a function of the random packet/event drop probability applied independently at every layer. C) Accuracy as a function of the standard deviation (SD) of random noise values added to each spike time in the network (spiking time jitter). On fine-tuning for 10 epochs, the performance drop due to quantization, jitter, and random packet drops can be significantly improved. The average accuracy across 10 trials is reported here. D) The test accuracy obtained using the $\pi^{2}_{s}$ algorithm on the CIFAR10 and CIFAR100 datasets. The accuracies for spiking sparsity ratios ($P/d$) 1 and 0.3 are reported and compared with the state-of-the-art [r18]. The sparsity ratio is set to 1 for the input layer for all the experiments.
@@
We perform similar experiments on the F-MNIST dataset trained on the LeNet-5 network architecture comprising 60K parameters. We train the traditional MAC-based network and achieve an accuracy of 90.47%. On porting the weights to an equivalent $\pi^{2}$-NN network with appropriately chosen values of $K$ and $\alpha$, we achieve an initial test accuracy of 85.23%. Re-training the $\pi^{2}$-NN improves the accuracy to 90.7%. Training a $\pi^{2}$-NN network from scratch with different random initializations leads to a 1% accuracy drop on average. For comparison, [stanojevic2024] reports a test accuracy of 90.94% using a TTFS (Time-To-First-Spike) spiking neural network (SNN) variant of LeNet, while [lenet] reports 90.1% using an SNN-based LeNet model trained on the FMNIST dataset.
@@
Nevertheless, upon further training using the initialized weights, the $\pi^{2}$-ResNet-9 achieves a competitive maximum accuracy of 94.23% on CIFAR-10. Prior works on ANN-based models report accuracies of 94.29% and 90.8% in [cifar103] and [cifar102], respectively. A custom 9-layer residual convolutional network, DavidNet, achieves 94.08% accuracy [cifar101], while a spiking ResNet-18 model (which has around 11.2M parameters) demonstrates 94.25% accuracy in [cifar104]. Notably, 94% is also considered the benchmark for human-level performance on CIFAR-10 [cifar105]. We also experimented with training the network naively from scratch, without weight porting and knowledge distillation strategies, and achieved an accuracy of 89%.
@@
For the CIFAR-100 dataset, the $\pi^{2}$-based ResNet-9 model achieves a maximum test accuracy of 75.5%. In comparison, previous studies report accuracies of 75.41% and 68.25% on ResNet-9-based DNN models, in [cifar1001] and [cifar1002], respectively. In [cifar1003], performance is further improved through model ensembling, resulting in an accuracy of 77%. However, the accuracy achieved using a single ResNet-9 model in that study is 72.21%. Further, researchers report an accuracy of 74.24 % on ResNet-18-based SNNs in [cifar104]. These comparative results are summarized in Fig. 6A. Training the $\pi^{2}$-NN naively, from scratch, without weight porting and knowledge distillation strategies, resulted in an accuracy of 69%. Further details of the training configuration are provided in the Methods section 4.3.
Training $\pi^{2}$-NNs on GPUs is both memory-intensive and computationally demanding due to their differential formulation and the nature of their core operations. While GPUs are highly optimized for parallel matrix–vector multiplications, $\pi^{2}$-NN training requires sorting and event-driven operations that are not inherently GPU-friendly [causal]. As a result, we were unable to demonstrate results for significantly larger network architectures. All our experiments were conducted on a single 40GB NVIDIA A100 GPU. Training a $\pi^{2}$-ResNet-9 network on the CIFAR datasets with a batch size of 16 required approximately 34 GB of GPU memory, thus constraining the training of larger-scale networks. It is important to note that the GPU experiments serve solely to verify the correctness of the learning rule and representational capacity of $\pi^{2}$-NN; the computational advantages emerge only when deployed on hardware where delay and sorting are physical rather than simulated operations. Thus, the results should not be interpreted as a demonstration of the computing in interconnects capability but rather as numerical emulations of the underlying dynamics of the time-domain equations.
@@
Fine-tuning for hardware: One critical architectural lever for optimizing hardware efficiency is the choice of the parameter K. Reducing the value of K improves computational sparsity (as fewer inputs need to be processed, as portrayed in 5C,D), reduces hardware complexity, and minimizes the buffer and queue size requirements. Specifically, a simplified processor design can be implemented by setting K=1 in Eq. 1. The $\pi^{2}$-Rx node generates an event immediately upon receiving the first input event, reducing the neuron model to a basic processing element—a switch. This kind of paradigm evaluates the interconnects’ computational capabilities alone. While this approach significantly improves throughput by minimizing computational complexity to that of a switching operation (Methods section 4.8), it comes at the cost of reduced accuracy (Fig. 6 B,C,D). Further, the $K=1$ networks also eliminate the memory access overhead of fetching the pre-computed values (membrane potential) from memory, thus enhancing the energy efficiency of $\pi^{2}$ networks [1time_step, 1time_step1]. However, it is important to note that the performance of $\pi^{2}$-NNs, as observed in software simulations, may degrade in real hardware due to physical limitations and non-idealities of the interconnection networks. Therefore, the network must be designed to be robust against hardware-level effects such as quantization, temporal noise (jitter), and packet drops.
A related and equally important consideration for maintaining higher energy efficiency in neuromorphic systems is spiking sparsity. In the baseline $\pi^{2}$ formulation, each neuron emits two differential spikes/neuron (Eq. 5-9), which increases event traffic. However, the fraction of spikes per neuron per input—the standard measure of communication sparsity—directly impacts energy consumption in neuromorphic systems [stanojevic2024]. Sparse spiking is beneficial because (i) synaptic weight memory reads dominate energy in digital neuromorphic implementations, and unused synapses incur no cost, and (ii) spike-transmission cost becomes the dominant energy term as network size grows [stanojevic2024]. Additionally, achieving and controlling high spiking sparsity is crucial for efficient bandwidth utilization.
@@
Analyzing the tradeoffs between K and computational accuracy: We evaluate the computational capabilities of the interconnects using the $K=1$ paradigm within a 3-layer fully connected network trained on the MNIST dataset. Initially, $K$ is fixed at 1 in the hidden layer, while the output layer is allowed to use $K>1$ during training. Subsequently, we constrain $K$ to 1 in the output layer as well, to assess the network’s performance under fully constrained conditions. The results are summarized in Fig. 6B. A baseline accuracy of 97.7% was previously reported for a 3-layer MLP with 100 hidden neurons [temp]. Applying the $K=1$ constraint in the hidden layer of the same architecture results in a 2% drop in accuracy. However, increasing the hidden layer dimension to 500 restores the original performance. The trade-off in accuracy when K is constrained to 1 in the output layer as well is also reported.
@@
We extend a similar experiment to the $\pi^{2}$-NN-based LeNet-5 architecture trained on the FMNIST dataset. As illustrated in Fig. 6C, imposing the $K=1$ constraint on the first four layers of the network leads to an approximate 10% reduction in test accuracy relative to the baseline. To mitigate this drop, we relax the constraint by setting $K=10$ across all layers. This results in a more modest accuracy degradation of approximately 3%, highlighting the trade-off between computational sparsity and model performance. We evaluated simulation latency across different K configurations for a deeper LeNet-5 (five-layer CNN) trained on the FMNIST dataset and observed that the latency values span a very similar range across the tested settings. Specifically, when the configurations of K across the 5 layers were set to be - [12,50,85,25,75], [1,1,1,25,75], [10,10,10,10,10], we observed a simulation latency in the range of 30-35 time units. Reducing K to 1 did not lead to a significant reduction in latency. These results indicate that, while smaller K values can offer latency benefits in shallow networks, such benefits diminish in deeper architectures. Thus, K should be viewed as a design parameter that trades computational sparsity for accuracy, rather than as a knob that guarantees latency reduction at arbitrary depth.